solve using elimination method If system has no solution or infinate state this
-12x+27y=60
-6x-6y=30

Answer by mananth(16949)

-12 x + 27 y = 60 .............1

-6 x -6 y = 30 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 9
-24 x 54 y = 120
-54.00 x -54 y = 270
Add the two equations
-78.00 x = 390
/ -78.00
x = -5.00
plug value of x in (1)
-12.00 x + 27.00 y = 60.00
60.00 + 27.00 y = 60.00
27.00 y = 60.00 -60.00
27.00 y = 0.00
y = 0.00